2. A square support unit in a television is made with a side measuring 3 centimeters. A new model being designed for the upcoming year will have a larger square with a side measuring 3.2 centimeters. By how much will the area of the square be increased?
Thank you for your help, it is greatly appreciated.
AD

Answer by checkley77(12844)

3^2=9
3.2^2=10.24
10.24/9=1.138 or the area is increased by 13.8%.